OpenStudy (anonymous):

Find the general solution of dy/dx+3y=e^3x.

OpenStudy (loser66):

question: is it not first order linear equation and we should find out the integrating factor?

zepdrix (zepdrix):

Oh this is first order :) derrrrp lol. Bah I'm a little tired I think.

OpenStudy (loser66):

@zepdrix I didn't study it yet, I don't know how to solve.

OpenStudy (loser66):

integral p(x) = 3x

OpenStudy (dumbcow):

for diff equ \[y' + p(x)y = q(x)\] integrating factor is: \[\large e^{\int\limits p(x) dx}\]

OpenStudy (loser66):

thank you

OpenStudy (loser66):

continue please, I save time,

OpenStudy (loser66):

ok, let me practice for next semester. so integrating factor is e^3x

OpenStudy (dumbcow):

yes

OpenStudy (loser66):

time both sides by e^3x to gety'e^3x + 3e^3xy= (e^3x)^2

OpenStudy (loser66):

yes or no?

OpenStudy (dumbcow):

correct

OpenStudy (loser66):

factor to get d/dt (e^3x*3y)=e^3x)^2

OpenStudy (dumbcow):

looks like you have it pretty figured out...why dont you finish it

OpenStudy (dumbcow):

oh you dont need that extra "3" --> d/dx(e^3x *y)

OpenStudy (loser66):

why? the original problem has 3y , so time integrating factor in , why 3 disappears?

OpenStudy (dumbcow):

because of chain rule derivative of e^3x = 3*e^3x

OpenStudy (dumbcow):

ok so we have \[\large e^{3x} y' +3e^{3x} y = e^{6x}\] left side is in form of product rule \[\large ->e^{3x}*y' + (e^{3x})' *y\] so you can say \[\large (e^{3x} y)' = e^{6x}\] now integrate both sides \[\large e^{3x} y = \frac{1}{6} e^{6x} +C\] solve for "y" \[\large y = \frac{1}{6}e^{3x} +C e^{-3x}\]

OpenStudy (loser66):

since we don't have initial condition, no need to solve for C. If we have y_0 we have to solve for C and plug back to this y , right?

OpenStudy (dumbcow):

where is @idealist

OpenStudy (dumbcow):

yes correct...its asking for general solution so constant is ok also they would have given initial conditions otherwise

OpenStudy (dumbcow):

haha ok and yw

OpenStudy (dumbcow):

right, its a wierd system they have now...a lot of 99's out there

OpenStudy (dumbcow):

:)